solve by factoring x squared +7=28-4x
Answer by kiru_khandelwal(79)

x^2 + 7 = 28 -4x
=> x^2 + 4x +7 - 28 = 0
=> x^2 + 4x - 21 = 0
=> x^2 + 7x - 3x - 21 = 0
=> x(x+7) -3(x+7) = 0
=> (x+7)(x-3) =0
=> x+7 = 0 or x-3 =0
=> x = -7, 3